an investor deposits $100,000 into two accounts. Twice the larger deposit is $25,000 less than three times the smaller deposit. Find the amount deposited in each account.
Answer by arunpaul(104)

Let account A have diposit of amount = x (larger)
account B have diposit of amount = y smaller)
so x+y = $100000 ----1
again as per question
2x = $25,000 + 3y
i.e 2x -3y = 25000 -----2
multiply equation 1 with 2 and subtract from equation 2
2x -3y = 25000
- 2x +2y = 200000
(subtracting we get)
-------------------------
- 5y = (25000-200000)
- 5y = - 175000
y = (175000/5) = 35000
so x = 100000-35000= 65000
so deposit A = 65000
and deposit B = 35000